UUCP Tips

Donnie Barnes, djb@redhat.com V1.0, Jan. 22, 1996

1 Preface

“Red Hat Tips” are documentation meant to help Red Hat users with specific tasks. Some of these documents are for new users, some are for advanced users. Hopefully each document will also be of help for both new and advanced users. If you have contributions to make, please send them to tech-sup@redhat.com. If you have changes that need to be made to individual Tips, send them to the author of that document.

2 Introduction

The text of this HOWTO was contributed by Julien Michielsen. If you have problems or questions with any of this, please subscribe to the redhat-list mailing list and send them there. For more info on that list, see

http://www.redhat.com.

For those of you also would like to setup uucp, my configuration files might be useful to you. This documentation covers the use of UUCP with smail instead of sendmail. Red Hat ships with sendmail configured, but you can uninstall it (rpm -u sendmail) and install the smail RPM from the contrib directory on your CD or ftp.redhat.com. If you want sendmail, you’ll have to figure out yourself how that works.

3 The Basics

I am a leaf-site in the hobby.nl domain. My provider is hgatenl.hobby.nl. A local second provider (murphy.hobby.nl) can also be called. He is my forwarder for news. I am still in the process of getting news setup. Therefore I have not tested the setup of murphy yet.

Things to do:

• make sure /dev/modem exists. If not, make a symbolic link between modem and the cua?-port your modem is on (cua0=com1, cua1=com2, ....).

   • permissions of my /dev/cua3 are crw-rw-rw-

• create the file /var/log/uucp, ownership uucp.uucp, permissions -rw-rw-r--

• create the directory /etc/uucp, and create the files dial, port and sys in there. Ownership /etc/uucp is uucp.uucp. (Also of the files in /etc/uucp.)

4. Sample Configuration Files 2

4 Sample Configuration Files

/etc/uucp/sys:
remote-send ~
remote-receive ~
local-send ~
local-receive ~
command-path /usr/sbin
commands rmail
#time Wk1800-0700,Sa,Su
time any
port ACU
max-remote-debug none
#debug chat
#protocol-parameter i
system murphy
commands rmail
call-login Ukennis
#          ^may be specific for my situation
call-password NotKnown
phone 0109876543
system hgatenl
commands rmail
call-password MySecret
chat "" \d\d\d ogin: \dSkennis assword: MySecret "" \d\d Uucp u
#                      ^^specific ??
phone 0206781234
/etc/uucp/dial:
dialer hayesnormal
chat "" ATZ\r\c\d\d OK\r ATQ0M1E0\r\c OK\r AT\&K1\r\c OK\r ATDT\T\r\c CONNECT \d\d\d\r
chat-fail RING\r
chat-fail NO\sCARRIER
chat-fail ERROR
chat-fail NO\sDIALTONE
chat-fail BUSY
chat-fail NO\sANSWER
chat-fail VOICE
complete \d\d+++\d\dATH\r\c
abort \d\d+++\d\dATH\r\c
/etc/uucp/port:

5. Smail Configuration 3

port ACU
type modem
baud 38400
device /dev/modem
carrier true
dialer hayesnormal

5 Smail Configuration

Subsequently install the smail RPM. I used the smail.3.1.29.i386.rpm, which worked fine. In the /etc/smail directory I made the following entry in /etc/smail/config :

hostname=kennis.hobby.nl
domains=hobby.nl
smart_path=hgatenl
# max_message_size=512k
# domains=foo.bar
# smart_path=isp
# uucp_name=hname.foo.bar
# visible_name=foo.bar
# auth_domains=foo.bar
# more_hostnames=nonfoo:nonfoo.foo.bar
Create /etc/smail/maps/map:

#N      foo.bar foo2.bar2
#S      AT 486/RedHat Linux 1.2.13
#O      organization
#C      contact
#E      administration (email)
#T      phone
#P      address
#R
#U      hosts connected via uucp
#W      created/edited by
#
# hname isp
# hname foo.bar
# hname = foo.bar
# hname = hname.foo.bar
kennis  .hobby.nl(LOCAL)
kennis = kennis.hobby.nl
hgatenl kennis(DIRECT)

6 You’re Done!

Hope I didn’t forget any adjustments I made, and that this will help to get you in the air. Best of luck!

7 Copyright Notice

This document is Copyright (C) 1996 by Red Hat Software. Redistribution of this document is permitted as long as the content remains completely intact and unchanged. In other words, you may reformat and reprint or redistribute only.